# for Safety Risks, Jailbreaks, and Refusals of LLMs

### v2

@@ 2 The Status Quo of Safety Moderation Tools for LLMs
To evaluate harm detection in adversarial prompts, we sample a uniform mixture of 250 benign and 250 harmful prompts from the validation set of WildJailbreak (WJ[20], a large-scale dataset consisting of adversarial and vanilla queries from diverse harm categories. To evaluate nuanced refusal detection, we use our novel benchmark XSTest-Resp, which we describe in §3.2. For comparison, we also show results on our third task, response harmfulness, on XSTest-Resp.
@@
We evaluate both open-source and closed tools on these tests. Among open-source tools we test four models instruction-tuned to identify harmfulness in both prompts and responses: Llama-Guard [16], Llama-Guard2 [26], Aegis-Guard-Defensive [13], and Aegis-Guard-Permissive [13]. For these models, to label refusal we mark a response as compliance if the output is harmful, and as a refusal if the output label is safe. We show results for harmful prompts only (RR(h) in Table 2), as this mapping is more likely to succeed in this setting. We also test LibrAI-LongFormer-ref, which is trained to detect refusals on the Do-Not-Answer benchmark [36] and used for evaluations in TrustLLM [32]. Lastly, we test a keyword-based classifier, which identifies refusals using a predefined list of refusal keywords (see our list of keywords in Appendix F.3). We evaluate the latter two models only on refusal detection. We also assess two API-based tools: GPT-4  [^1] We use gpt-4-0125-preview for all our experiments. [1] and OpenAI Moderation API [28]. For GPT-4 we experiment with various prompts and select the one that results in best performance. We provide this prompt in Table 15 of Appendix F.2. OpenAI Moderation API returns labels for toxicity, which we use for prompt and response harmfulness.
@@ 3.1.1 Prompt Construction
Vanilla harmful synthetic prompts. We synthetically generate harmful prompts covering a broad range of risk scenarios. Inspired by Weidinger et al. [37], we consider 13 harm subcategories in four high-level categories: privacy, misinformation, harmful language, and malicious uses. Table 10 in Appendix C.6 lists harm categories with statistics of our data. To generate targeted, realistic, and diverse harmful scenarios, we design a structured generation pipeline, detailed in Appendix C.1.
Vanilla benign synthetic prompts. Similarly to WildJailbreak [20], to enable maximally precise detection of prompt harmfulness, we additionally include two types of benign contrastive prompts: 1) benign prompts that superficially resemble unsafe prompts, motivated by the 10 exaggerated categories [^2] Homonyms, figurative language, safe targets, safe contexts, definitions, real discrimination/nonsense group, nonsense discrimination/real group, historical events, public privacy, and fictional privacy. from XSTest [30], and 2) benign prompts discussing sensitive but safe topics. These vanilla benign contrastive prompts are generated with GPT-4. Refer to Table 9 in Appendix C.2 for the GPT-4 prompt used to generate the queries and to Table 7 for examples.
Adversarial benign and harmful synthetic prompts. We employ the WildTeaming [20] framework to convert our generated vanilla prompts—harmful and benign—to adversarial prompts. This framework mines jailbreaking tactics from in-the-wild user-LLM interactions, and composes selections of tactics to transform vanilla requests into diverse adversarial counterparts.
@@ 3.1.2 Compliance and Refusal Construction
GPT-4 complex response generations. We additionally use GPT-4 to generate items capturing a set of targeted response types in observed challenge areas for refusal classification. We conduct error analysis on a classifier trained on an early prototype of WGTrain, and identify several categories of nuanced responses that the classifier mislabels. The majority of these categories are compliances containing caveats, warnings, etc. We then construct one-shot prompts for GPT-4 to generate synthetic responses with similar properties to enhance coverage of complex responses. [^4] Most items take user prompts by XSTest category from the vanilla benign prompt set, while for a few categories the user prompts were also synthetically generated. See Appendix C.5 for all categories and prompts.
@@ 3.1.3 Filtering, Auditing, and Sampling
We sample from our other data sources (on which filtering/audit is not needed) as follows: from complex response, 1,167 each of refusal, compliance, and prompt-only items; from In-the-Wild data, 944 harmful prompts and 944 benign prompts; and from annotator-written data, 7,361 benign and 2,130 harmful prompts. [^5] See Appendix C.7 for further breakdowns. In-the-Wild and annotator-written items contain only prompts.
@@ 3.2 WildGuardTest: A High-Quality Human-Annotated Test Moderation Dataset
To construct WGTest, we start with a test split of 1,725 prompt-response pairs from the synthetic vanilla and adversarial data as described in §3.1. We then collect annotations from three independent annotators for each prompt-response pair on prompt harmfulness, response refusal, and response harmfulness. Fleiss Kappa [10] scores are 0.55, 0.72, and 0.50 for the three tasks, indicating moderate to substantial agreement. We apply majority voting to obtain gold labels, removing items that fail to reach at least two-way agreement. [^6] Lack of agreement is possible because we include an “unsure” answer choice. Finally, we run a prompted GPT-4 classifier on the dataset and manually inspect items on which the output mismatches the chosen annotator label, to further audit the ground-truth labels. See Table 10 in Appendix C.5 for the WGTest label breakdowns and harm category splits according to our risk taxonomy, and Appendix G for more annotation details.
@@ 3.3 Training WildGuard with WildGuardTrain
With WGTrain, we instruction-tune WildGuard using Mistral-7b-v0.3  [19] as a base model. [^7] We experiment with various base models (see Appendix H.1 for comparisons). Mistral-7b-v0.3 shows top performance, but difference margins are small, emphasizing primary importance of WildGuardTrain. We design a unified input and output format to capture the three tasks, providing the user prompt and model response as inputs, and training the model to output elements corresponding to all three tasks. See more training details in Appendix D.
@@ 4.1 Evaluation Setups
Evaluation benchmarks. We test WildGuard and relevant baselines on ten publicly-available safety benchmarks, as well as our WGTest across all three tasks. For public benchmarks on prompt harmfulness, we use ToxicChat [23], OpenAI Moderation [24], AegisSafetyTest [13], SimpleSafetyTests [35], and HarmBenchPrompt [25]. For public benchmarks on response harmfulness, we use HarmBenchResponse [25], SafeRLHF [9], BeaverTails [18], and XSTest-Resp  [^8] We count XSTest-Resp among public benchmarks because the prompts originate from XSTest [30].. The only evaluation for refusal detection other than WGTest is XSTest-Resp, which we discuss in §2. Table 14 in Appendix E shows statistics of each benchmark. We report F1 scores for all evaluations.
Existing Safety Moderation Models. As in Section 2, we test four LM-based moderation tools trained for prompt and response harmfulness detection: Llama-Guard, Llama-Guard2, Aegis-Guard-Defensive and Aegis-Guard-Permissive [16, 26, 13]. We test five additional models for response harmfulness classification: BeaverDam [18], LibrAI-LongFormer-harm [36], MD-Judge-v0.1  [22], and two Harmbench classifiers from Mazeika et al. [25] (HarmBench-Llama, HarmBench-Mistral). For refusal detection, we test LibrAI-LongFormer-ref [36]. See Appendix §F for additional model details. For closed tools, as in Section 2, we test the OpenAI Moderation API and GPT-4.
@@ Acknowledgements
Appendix A Limitations
As is standard and generally unavoidable in safety research, we make commitments about what specific categories of content constitute harmful categories. These may differ from the categories that others may prefer. The risk categories we focus on are listed in Table 10 of the Appendix. We also acknowledge that our risk taxonomy might not cover all potential harms, but we provide the risk categories that existing datasets/benchmarks cover in Appendix E, and our risk taxonomy and the specific pinpoint topics WildGuardMix addresses cover these categories as well, leading to strong performances on these existing benchmarks.
Along similar lines, it is necessary that we make commitments on definitions of refusal, which may not perfectly match others’ preferences. The definition of refusal is shown in Figure 4, displaying the instruction of what makes a model response a refusal. Our definition of refusals includes different types of scenarios, such as “redirecting refusal” and “selective refusal”, along with the examples of compliances that can be confused with refusals, such as “refusal then compliance” and “correction compliance”. Addressing these complex behaviors remains a challenge, and we look forward to continuing to refine our approaches in future work.
Finally, an area that we have omitted in WildGuard is finer-grained classification of harm category. Though the three-task setup of WildGuard is already broad in coverage, this additional capability will be an interesting direction to explore in future work.
Appendix B Ethical Considerations
Appendix C More Details about WildGuardMix
C.1 A Pipeline for Vanilla Harmful Synthetic Prompts
To enable creation of more targeted, realistic, and diverse harmful scenarios, we design the following pipeline: 1) We first identify topic-specific dimensions in each risk category (e.g., harmful language) and subcategory (e.g., stereotypes) with detailed descriptions of the subcategories. Then we manually list pinpoint topics (e.g., body shaming, skin discrimination, or immigrant status discrimination) within these subcategories, enhancing GPT-4’s ability to generate more targeted and varied scenarios. 2) After pinpoint topics are identified, we provide GPT-4 with these topics along with the risk category, subcategory, and definition of the subcategory Optionally, we use few-shot examples to enhance the diversity of the generated prompts. After generating the prompts, we run a minhash lexical deduplication based on 3-gram overlap [27], applying a threshold of 0.5.
C.2 GPT-4 Instructions for Generating Vanilla Prompts
@@
C.3 Examples for Vanilla Benign Prompts
@@
C.4 Language Model Instructions for Generating Compliances and Refusals
@@
C.5 Complex Response Generation
@@
C.6 Fine-grained Risk Taxonomy of WildGuardMix
@@
C.7 Public Annotator-written Data
@@ C.8 Examples for WildGuardMix
Appendix D WildGuard Training Details
@@
Appendix E Public Benchmarks for Evaluations
@@
Appendix F Additional Details about Existing Moderation Tools
F.1 Open-source LM-based Tools
@@ F.2 GPT-4 Classification
F.3 Keyword-based Refusal Detection
@@
Appendix G Annotation Details
@@
Appendix H Extended Results
H.1 Base Model Comparison for WildGuard
@@
H.2 Full Evaluation Results
@@
H.3 More Results for WildGuard Demonstrations as a Moderator in Human-LLM Interactions

### v3

@@ 2 The Status Quo of Safety Moderation Tools for LLMs
To evaluate harm detection in adversarial prompts, we sample a uniform mixture of 250 benign and 250 harmful prompts from the validation set of WildJailbreak (WJ[20], a large-scale dataset consisting of adversarial and vanilla queries from diverse harm categories. To evaluate nuanced refusal detection, we use our novel benchmark XSTest-Resp, which we describe in §3.2. For comparison, we also show results on our third task, response harmfulness, using XSTest-Resp.
@@
We evaluate both open-source and closed tools on these tests. Among open-source tools we test four models instruction-tuned to identify harmfulness in both prompts and responses: Llama-Guard [16], Llama-Guard2 [26], Aegis-Guard-Defensive [13], and Aegis-Guard-Permissive [13]. For these models, to label refusal we mark a response as compliance if the output is harmful, and as a refusal if the output label is safe. We show results for harmful prompts only (RR(h) in Table 2), as this mapping is more likely to succeed in this setting. We also test LibrAI-LongFormer-ref, which is trained to detect refusals on the Do-Not-Answer benchmark [38] and used for evaluations in TrustLLM [32]. Lastly, we test a keyword-based classifier, which identifies refusals using a predefined list of refusal keywords (see our list of keywords in Appendix D.3). We evaluate the latter two models only on refusal detection. We also assess two API-based tools: GPT-4  [^1] We use gpt-4-0125-preview for all our experiments. [1] and OpenAI Moderation API [28]. For GPT-4 we experiment with various prompts and select the one that results in best performance. We provide this prompt in Table 15 of Appendix D.2. OpenAI Moderation API returns labels for toxicity, which we use for prompt and response harmfulness.
@@ 3.1.1 Prompt Construction
Vanilla harmful synthetic prompts. We synthetically generate harmful prompts covering a broad range of risk scenarios. Inspired by Weidinger et al. [39], we consider 13 harm subcategories in four high-level categories: privacy, misinformation, harmful language, and malicious uses. Table 10 in Appendix A.6 lists harm categories with statistics of our data. To generate targeted, realistic, and diverse harmful scenarios, we design a structured generation pipeline, detailed in Appendix A.1.
Vanilla benign synthetic prompts. Similarly to WildJailbreak [20], to enable maximally precise detection of prompt harmfulness, we additionally include two types of benign contrastive prompts: 1) benign prompts that superficially resemble unsafe prompts, motivated by the 10 exaggerated categories [^2] Homonyms, figurative language, safe targets, safe contexts, definitions, real discrimination/nonsense group, nonsense discrimination/real group, historical events, public privacy, and fictional privacy. from XSTest [30], and 2) benign prompts discussing sensitive but safe topics. These vanilla benign contrastive prompts are generated with GPT-4. Refer to Table 9 in Appendix A.2 for the GPT-4 prompt used to generate the queries and to Table 7 for examples.
Adversarial benign and harmful synthetic prompts. We employ the WildTeaming [20] framework to convert our generated vanilla prompts—harmful and benign—to adversarial prompts. This framework mines jailbreaking tactics from in-the-wild user-LLM interactions, and composes selections of tactics to transform vanilla requests into diverse adversarial counterparts. Specifically, WildTeaming first uses the OpenAI Moderation API to find potentially harmful user queries from LMSYS-Chat-1M and WildChat. Then we use prompted GPT-4 to decompose intent and jailbreak tactics from the queries. This mining process ends up with a pool for the jailbreak tactics; then we randomly sample 2-7 In-the-Wild tactics from the pool, and finally use the sampled tactics to transform vanilla prompts into adversarial prompts.
@@ 3.1.2 Compliance and Refusal Construction
GPT-4 complex response generations. We additionally use GPT-4 to generate items capturing a set of targeted response types in observed challenge areas for refusal classification. We conduct error analysis on a classifier trained on an early prototype of WGTrain, and identify several categories of nuanced responses that the classifier mislabels. The majority of these categories are compliances containing caveats, warnings, etc. We then construct one-shot prompts for GPT-4 to generate synthetic responses with similar properties to enhance coverage of complex responses. [^4] Most items take user prompts by XSTest category from the vanilla benign prompt set, while for a few categories the user prompts were also synthetically generated. See Appendix A.5 for all categories and prompts.
@@ 3.1.3 Filtering, Auditing, and Sampling
We sample from our other data sources (on which filtering/audit is not needed) as follows: from complex response, 1,167 each of refusal, compliance, and prompt-only items; from In-the-Wild data, 944 harmful prompts and 944 benign prompts; and from annotator-written data, 7,361 benign and 2,130 harmful prompts. [^5] See Appendix A.7 for further breakdowns. In-the-Wild and annotator-written items contain only prompts.
@@ 3.2 WildGuardTest: A High-Quality Human-Annotated Test Moderation Dataset
To construct WGTest, we start with a test split of 1,725 prompt-response pairs from the synthetic vanilla and adversarial data as described in §3.1. We then collect annotations from three independent annotators for each prompt-response pair on prompt harmfulness, response refusal, and response harmfulness. Fleiss Kappa [10] scores are 0.55, 0.72, and 0.50 for the three tasks, indicating moderate to substantial agreement. We apply majority voting to obtain gold labels, removing items that fail to reach at least two-way agreement. [^6] Lack of agreement is possible because we include an “unsure” answer choice. Finally, we run a prompted GPT-4 classifier on the dataset and manually inspect items on which the output mismatches the chosen annotator label, to further audit the ground-truth labels. See Table 10 in Appendix A.5 for the WGTest label breakdowns and harm category splits according to our risk taxonomy, and Appendix E for more annotation details.
@@ 3.3 Training WildGuard with WildGuardTrain
With WGTrain, we instruction-tune WildGuard using Mistral-7b-v0.3  [19] as a base model. [^7] We experiment with various base models (see Appendix F.1 for comparisons). Mistral-7b-v0.3 shows top performance, but difference margins are small, emphasizing primary importance of WildGuardTrain. We design a unified input and output format to capture the three tasks, providing the user prompt and model response as inputs, and training the model to output elements corresponding to all three tasks. See more training details in Appendix B.
@@ 4.1 Evaluation Setups
Evaluation benchmarks. We test WildGuard and relevant baselines on ten publicly-available safety benchmarks, as well as our WGTest across all three tasks. For public benchmarks on prompt harmfulness, we use ToxicChat [23], OpenAI Moderation [24], AegisSafetyTest [13], SimpleSafetyTests [36], and HarmBenchPrompt [25]. For public benchmarks on response harmfulness, we use HarmBenchResponse [25], SafeRLHF [9], BeaverTails [18], and XSTest-Resp  [^8] We count XSTest-Resp among public benchmarks because the prompts originate from XSTest [30].. The only evaluation for refusal detection other than WGTest is XSTest-Resp, which we discuss in §2. Table 14 in Appendix C shows statistics of each benchmark. We report F1 scores for all evaluations.
Existing Safety Moderation Models. As in Section 2, we test four LM-based moderation tools trained for prompt and response harmfulness detection: Llama-Guard, Llama-Guard2, Aegis-Guard-Defensive and Aegis-Guard-Permissive [16, 26, 13]. We test five additional models for response harmfulness classification: BeaverDam [18], LibrAI-LongFormer-harm [38], MD-Judge-v0.1  [22], and two Harmbench classifiers from Mazeika et al. [25] (HarmBench-Llama, HarmBench-Mistral). For refusal detection, we test LibrAI-LongFormer-ref [38]. See Appendix §D for additional model details. For closed tools, as in Section 2, we test the OpenAI Moderation API and GPT-4.
@@ 5 Related Works
Model Safety Taxonomy. Multiple previous works develop taxonomies to identify and break down potential model safety risks into categories. We draw inspiration for the taxonomy used for WildGuardMix from Weidinger et al. [39], which defines several broad risk areas with detailed sub-categories. Other works that define model safety taxonomies include Tedeschi et al. [34] and Vidgen et al. [37]. The taxonomies overlap significantly, but feature differences in the types of content explicitly or implicitly covered and the specificity of categorization within each broad risk area.
@@
6 Limitations
As is standard and generally unavoidable in safety research, we make commitments about what specific categories of content constitute harmful categories. These may differ from the categories that others may prefer. Since other models may be developed with different underlying harm taxonomies, they might demonstrate lower performance when evaluated on WildGuardTest because of discrepancies between our definition of harm and the model developers’ goals. The risk categories we focus on are listed in Table 10 of the Appendix. We also acknowledge that our risk taxonomy might not cover all potential harms, but we provide the risk categories that existing datasets/benchmarks cover in Appendix C, and our risk taxonomy and the specific pinpoint topics WildGuardMix addresses cover these categories as well, leading to strong performances on these existing benchmarks.
Along similar lines, it is necessary that we make commitments on definitions of refusal, which may not perfectly match others’ preferences. The definition of refusal is shown in Figure 4, displaying the annotator instruction for what makes a model response a refusal. Our definition of refusals includes multiple scenarios, such as “redirecting refusal” and “selective refusal”, along with the examples of compliances that can be confused with refusals, such as “refusal then compliance” and “correction compliance”. Addressing these complex behaviors remains a challenge, and we look forward to continuing to refine our approach in future work.
Finally, an area that we have omitted in WildGuard is finer-grained classification of harm category. Though the three-task setup of WildGuard is already broad in coverage, this additional capability is an interesting direction to explore in future work.
7 Ethical Considerations
@@ Acknowledgements
Appendix A More Details about WildGuardMix
A.1 A Pipeline for Vanilla Harmful Synthetic Prompts
To enable creation of more targeted, realistic, and diverse harmful scenarios, we design the following pipeline: 1) We first identify topic-specific dimensions in each risk category (e.g., harmful language) and subcategory (e.g., stereotypes) with detailed descriptions of the subcategories. Then we manually list pinpoint topics (e.g., body shaming, skin discrimination, or immigrant status discrimination) within these subcategories, enhancing GPT-4’s ability to generate more targeted and varied scenarios. 2) After pinpoint topics are identified, we provide GPT-4 with these topics along with the risk category, subcategory, and definition of the subcategory. Optionally, we use few-shot examples to enhance the diversity of the generated prompts. After generating the prompts, we run a minhash lexical deduplication based on 3-gram overlap [27], applying a threshold of 0.5.
A.2 GPT-4 Instructions for Generating Vanilla Prompts
@@
A.3 Examples for Vanilla Benign Prompts
@@
A.4 Language Model Instructions for Generating Compliances and Refusals
@@
A.5 Complex Response Generation
@@
A.6 Fine-grained Risk Taxonomy of WildGuardMix
@@
A.7 Public Annotator-written Data
@@ A.8 Examples for WildGuardMix
Appendix B WildGuard Training Details
@@
Appendix C Public Benchmarks for Evaluations
@@
Appendix D Additional Details about Existing Moderation Tools
D.1 Open-source LM-based Tools
@@ D.2 GPT-4 Classification
D.3 Keyword-based Refusal Detection
@@
Appendix E Annotation Details
@@
Appendix F Extended Results
F.1 Base Model Comparison for WildGuard
@@
F.2 Full Evaluation Results
@@
F.3 More Results for WildGuard Demonstrations as a Moderator in Human-LLM Interactions